### v1

# Gradient-Direction-Aware Density Control for 3D Gaussian Splatting

Abstract
The emergence of 3D Gaussian Splatting (3DGS) has significantly advanced novel view synthesis through explicit scene representation, enabling real-time photorealistic rendering. However, existing approaches manifest two critical limitations in complex scenarios: (1) Over-reconstruction occurs when persistent large Gaussians cannot meet adaptive splitting thresholds during density control. This is exacerbated by conflicting gradient directions that prevent effective splitting of these Gaussians; (2) Over-densification of Gaussians occurs in regions with aligned gradient aggregation, leading to redundant component proliferation. This redundancy significantly increases memory overhead due to unnecessary data retention. We present Gradient-Direction-Aware Gaussian Splatting (GDAGS), a gradient-direction-aware adaptive density control framework to address these challenges. Our key innovations: the gradient coherence ratio (GCR), computed through normalized gradient vector norms, which explicitly discriminates Gaussians with concordant versus conflicting gradient directions; and a nonlinear dynamic weighting mechanism leverages the GCR to enable gradient-direction-aware density control. Specifically, GDAGS prioritizes conflicting-gradient Gaussians during splitting operations to enhance geometric details while suppressing redundant concordant-direction Gaussians. Conversely, in cloning processes, GDAGS promotes concordant-direction Gaussian densification for structural completion while preventing conflicting-direction Gaussian overpopulation. Comprehensive evaluations across diverse real-world benchmarks demonstrate that GDAGS achieves superior rendering quality while effectively mitigating over-reconstruction, suppressing over-densification, and constructing compact scene representations with 50% reduced memory consumption through optimized Gaussians utilization. Our code can be accessed in the code repository https://github.com/zzcqz/GDAGS.
1 Introduction
Figure 1: Visualization of rendering performance and efficiency, with rendering frame rate and scene Gaussian count displayed in the upper left corner. The green box represents the optimized area, and the red box represents the poorly optimized area. Both GDAGS and AbsGS effectively densify and eliminate large Gaussians in over-reconstructed areas to restore fine details. However, AbsGS has too many over-densification Gaussians, which only contribute slightly to rendering quality. In contrast, GDAGS has a more compact and reasonable spatial distribution and only requires half of the Gaussians in Pixel-GS.
Novel View Synthesis (NVS) constitutes a foundational challenge in computer graphics and vision, aiming to reconstruct 3D scenes from monocular images to generate photorealistic novel perspectives. While neural implicit representations, such as Neural Radiance Fields (NeRF) mildenhall2021nerf , have significantly advanced NVS by modeling radiation fields through multilayer perceptrons (MLPs) and volumetric rendering, these methods struggle to reconcile rendering quality verbin2022ref with computational efficiency fridovich2022plenoxels . Emerging as a novel paradigm, 3D Gaussian Splatting kerbl20233d explicitly parameterizes scenes using adaptive Gaussian primitives. By integrating splat-based rasterization with adaptive density control strategies, 3DGS achieves real-time rendering of unprecedented fidelity while maintaining compact spatial representation.
However, the adaptive density control mechanism in 3DGS exhibits inherent limitations. Specifically, 3DGS’s densification criterion solely relies on the norm of view-space positional gradients, thereby overlooking the directional influence of gradients on the norm computation. During view-space positional gradient calculation for each Gaussian, the superposition of sub-gradient components with consistent directions amplifies the gradient norm. Conversely, counter-directional sub-gradient components mutually attenuate, resulting in reduced gradient norms. These directional conflicts prevent some Gaussians from effective densification, causing over-reconstruction phenomenon and perceptible local blurring. Meanwhile, Gaussians maintaining consistent gradient directions exhibit persistent densification tendencies, causing over-densification phenomena that generate redundant Gaussian primitives and significantly increase memory consumption. Recent work like AbsGS ye2024absgs has proposed using absolute subgradient values to enforce positive gradients for all Gaussians, thereby avoiding subgradient directional conflicts and partially mitigating the over-reconstruction issue. This approach further exacerbates the over-densification phenomenon, as shown in Figure 1.
To address the aforementioned dual challenges, we propose a gradient-direction-aware adaptive density control framework (GDAGS). This approach combines directional subgradient information and gradient magnitude data to establish a novel decision metric. The metric enables controlled densification of Gaussians with conflicting directional gradients while preventing redundant densification in directionally aligned cases. First, GDAGS calculates a normalized GCR to measure the directional consistency of subgradients for each Gaussian. The GCR value approaching 1 indicates highly consistent subgradient directions, whereas a value approaching 0 signifies substantial directional conflict. Subsequently, GDAGS employs GCR to parameterize a nonlinear dynamic weight $w_{i}$, which independently weights the view-space positional gradients of each Gaussian to form a new decision metric controlling densification behavior. During splitting operations, Gaussians exhibiting directional conflicts receive higher weights and are prioritized for splitting, while directionally consistent Gaussians obtain lower weights and are suppressed. Finally, GDAGS introduces a hyperparameter $p$ to explicitly control the number of Gaussians participating in densification, thereby achieving an indirect balance between performance and efficiency. We summarize our contributions as follows:
  • We utilize GCR as a gradient direction consistency measure for adaptive density control to identify Gaussians with consistent and divergent directions.
  • We propose a nonlinear dynamic weighting mechanism to independently weight the Gaussians in the scene, promoting correct densification and suppressing abnormal densification of the Gaussians.
  • Evaluated on three real-world datasets spanning 13 scenes, our method demonstrated superior visual quality while having a compact spatial representation that requires only 50% Gaussians and memory consumption.
2 Related work
2.1 Novel view synthesis
Novel View Synthesis has evolved through two interconnected trajectories: advancements in scene representation paradigms and optimizations of rendering efficiency. The field was revolutionized by NeRF mildenhall2021nerf , which introduced MLP-based implicit neural representations to map spatial coordinates to radiance and density values, achieving photorealistic results through volumetric rendering. However, NeRF’s computational demands spurred research bifurcating into quality refinement and efficiency optimization. Approaches like Mip-NeRF barron2021mip enhanced geometric stability using conical ray encoding and distortion regularization, while DVGO sun2022direct and Instant-NGP muller2022instant leveraged explicit 3D grids and hash encoding to achieve near-real-time rendering. This progression culminated in 3D Gaussian Splatting kerbl20233d , which redefined the paradigm by combining explicit Gaussian primitives with differentiable rasterization. By parameterizing scenes as adaptive 3D Gaussians, 3DGS bridged the gap between implicit continuity and explicit computational efficiency, achieving rendering speeds two orders of magnitude faster than NeRF while maintaining visual parity with state-of-the-art methods like Mip-NeRF360 barron2022mip .
2.2 3D gaussian splatting
The rapid adoption of 3DGS has driven innovations addressing its inherent limitations. Early improvements focused on geometric fidelity: Mip-Splatting yu2024mip integrated multi-scale primitives and low-pass filtering to mitigate aliasing, while GOF yu2024gaussian introduced composite gradient metrics to optimize Gaussian distributions. Subsequent methods explored structural constraints, such as 2DGSs huang20242d planar surface projections and RadeGSs zhang2024rade normal map-guided refinement. Frequency-aware approaches like FreGS zhang2024fregs employed spectral regularization to enhance high-frequency detail recovery, whereas HoGS liuHoGS utilized homogeneous coordinate reparameterization to improve distant object reconstruction. Despite these advances, critical challenges persist in adaptive density control. ABSGS ye2024absgs enforced gradient direction uniformity to resolve over-reconstruction but indiscriminately amplified outliers, while PixelGS zhang2024pixelgs traded memory efficiency for geometric precision by weighting gradients via pixel coverage—a strategy incurring prohibitive storage costs and frame rate penalties. Our work addresses these unresolved trade-offs through directional gradient awareness, systematically balancing precision, efficiency, and memory constraints.
3 Method
Section 3.1 establishes the methodological motivation for GDAGS, and Section 3.2 outlines 3DGS fundamentals. Sections 3.3.1 and 3.3.2 detail GDAGS’ core components, with full algorithmic implementation provided in Appendix A. The complete pipeline is shown in Figure 2.
Figure 2: The pipeline of GDAGS. First, for each Gaussian, GDAGS computes the GCR to quantify the directional coherence of its subgradients. Subsequently, this GCR metric is mapped through a nonlinear dynamic weighting function to generate per-Gaussian gradient weights, which modulate the view-space positional gradient magnitudes and produce a refined decision metric. Finally, this decision metric is compared against a predefined threshold to dynamically regulate densification. The detailed algorithm pseudocode is presented in Appendix A.
3.1 Motivation
The density control mechanism in 3D Gaussian Splatting fundamentally relies on comparing view-space positional gradient norms with predefined thresholds. However, the directional characteristics of subgradients implicitly modulate these gradient norms, either attenuating or amplifying the magnitude. This manifests as undesirable phenomena, including local blurring and excessive Gaussian proliferation, ultimately degrading rendering quality. To address these limitations, we propose to explicitly incorporate subgradient directional information into a nonlinear weighting function to construct a novel decision metric for resolving over-reconstruction and over-densification issues. Specifically, during splitting operations (targeting larger Gaussians), the GCR is utilized to determine the weighting strategy: Gaussians with smaller GCR values (indicating directional conflicts) receive amplified decision metrics through nonlinear weighting, thereby prioritizing their densification. Conversely, Gaussians with larger GCR values (exhibiting directional consistency) obtain diminished decision metrics and are suppressed. In contrast, during cloning operations (targeting smaller Gaussians), the weighting mechanism is inverted: low-GCR Gaussians are suppressed while high-GCR Gaussians are promoted for densification. This is because the Gaussian clones of high GCR can smoothly move towards the gradient direction, while the Gaussian clones of low-GCR stack locally. By systematically integrating both gradient directional information and magnitude characteristics into the decision metric, GDAGS optimizes density control mechanisms. This approach not only mitigates the inherent limitations of conventional 3DGS but also significantly reduces both the required number of Gaussians and memory consumption. Additional technical details are provided in Appendix B.
3.2 Preliminary
3DGS represents scenes using a set of learnable anisotropic 3D Gaussian primitives. Each Gaussian is parameterized by a position $\mu\in\mathbb{R}^{3}$, an opacity $\alpha\in[0,1]$, a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, and spherical harmonics (SH) arfken2011mathematical coefficients for view-dependent color. The covariance matrix $\Sigma$ is decomposed into a rotation matrix $R$ and a scaling matrix $S$ to ensure positive semi-definiteness:
$\Sigma=RSS^{\top}R^{\top},$ (1)
where $S$ is a diagonal matrix parameterized by a vector $s\in(\mathbb{R}_{+})^{3}$, and $R$ is a rotation matrix parameterized by a quaternion $q\in\mathbb{R}^{4}$. The Gaussian distribution is defined as:
$G(\mathbf{x})=\exp\left(-\frac{1}{2}(\mathbf{x}-\mu)^{\top}\Sigma^{-1}(\mathbf{x}-\mu)\right),$ (2)
where $\mathbf{x}\in\mathbb{R}^{3}$ is an arbitrary position.
During rendering, 3D Gaussians are projected into 2D camera space as $G\mathbf{{}^{\prime}}$, with pixel color $C(\mathbf{x}^{\prime})$ computed via alpha-blending:
$C(\mathbf{x}^{\prime})=\sum_{k=1}^{K}c_{k}\alpha_{k}G^{\prime}k(\mathbf{x}^{\prime})T_{k},\quad T_{k}=\prod{j=1}^{k-1}\left(1-\alpha_{j}G^{\prime}_{j}(\mathbf{x}^{\prime})\right),$ (3)
Here, $K$ denotes the count of Gaussians influencing the pixel, $c_{k}$ represents the view-dependent SH-based color, and $\alpha_{k}$ combines opacity with the Gaussian’s spatial contribution at $\mathbf{x}^{\prime}$.
To address sparse input from structure-from-motion (SfM) schonberger2016structure , a densification strategy is employed. Under-reconstructed regions (missing geometric features) are addressed by cloning Gaussians, increasing their number and volume. Over-reconstructed regions (large Gaussians covering small areas) are split into smaller Gaussians, maintaining volume but increasing count. Densification decisions are based on view-space gradients. For a Gaussian $G_{i}$ with projection $\mu_{i}^{k}$ under viewpoint $k$ and loss $L_{k}$, the average gradient magnitude is:
$\nabla_{\mu_{i}}L=\frac{1}{M}\sum_{k=1}^{M}\sqrt{\left(\frac{\partial L_{k}}{\partial\mu_{i,x}^{k}}\right)^{2}+\left(\frac{\partial L_{k}}{\partial\mu_{i,y}^{k}}\right)^{2}},$ (4)
where $M$ is the number of viewpoints involved in training. Split Gaussians into smaller components if $\nabla_{\mu_{i}}L>\tau_{p}$ and $\Sigma_{3D}^{i}>\tau_{s}$. Clone Gaussians to preserve local density if $\nabla_{\mu_{i}}L>\tau_{p}$ and $\Sigma_{3D}^{i}\leq\tau_{s}$.
3.3 GDAGS
3.3.1 Directional consistency measurement
For each Gaussian observed across $V$ views, we compute directional consistency metric GCR as $\mathcal{C}_{i}$:
$\mathcal{C}_{i}=\frac{\|\sum_{pixel}\nabla_{i,pixel}^{v}\|_{2}}{\sum_{pixel}\|\nabla_{i,pixel}^{v}\|_{2}+\epsilon},$ (5)
where pixelpixel denotes a pixel in the 2D space, and $\nabla_{i,pixel}^{v}\in\mathbb{R}^{2}$ represents the subgradient component projected onto each pixel in view $v$. According to the Cauchy-Schwarz Inequality $0\leq\|\sum_{pixel}\nabla_{i,pixel}^{v}\|_{2}\leq\sum_{pixel}\|\nabla_{i,pixel}^{v}\|_{2}$, it can be seen that the value of $\mathcal{C}_{i}$ is strictly controlled between [0,1], and GCR isolates the influence of gradient amplitude and only represents the consistency of gradient direction. This formulation captures:
  • High Consistency ($\mathcal{C}_{i}\to 1$): Gradient directions are consistent, their magnitudes effectively accumulate.
  • Low Consistency ($\mathcal{C}_{i}\to 0$): Conflicting gradient directions result in mutual cancellation of magnitudes.
The denominator $\sum_{pixel}\|\nabla_{i,pixel}^{v}\|_{2}$ captures the total gradient activity agnostic to directional variations, while the numerator $\|\sum_{pixel}\nabla_{i,pixel}^{v}\|_{2}$ quantifies the net directional alignment.
3.3.2 Nonlinear dynamic weight design
Let $\theta_{i,v}\in[0,2\pi)$ denote the gradient direction of Gaussian $i$ in view $v$, modeled as samples from a von Mises-Fisher (vMF) distribution:
$p(\theta_{i,v}|m_{i},\kappa_{i})=\frac{\kappa_{i}}{2\pi I_{0}(\kappa_{i})}e^{\kappa_{i}\cos(\theta_{i,v}-m_{i})},$ (6)
where $m_{i}$ is the mean direction and $\kappa_{i}>0$ the concentration parameter. The empirical directional consistency $\mathcal{C}_{i}$ relates to $\kappa_{i}$ through:
$\mathcal{C}_{i}=\frac{I_{1}(\kappa_{i})}{I_{0}(\kappa_{i})}\approx 1-\frac{1}{2\kappa_{i}}+\mathcal{O}(\kappa_{i}^{-2}),$ (7)
where $I_{n}$ denotes the modified Bessel function of order $n$.
Then, we define the target distribution $q^{*}(\nabla)$ that emphasizes directionally inconsistent regions:
$q^{*}(\nabla)=\frac{1}{Z}\|\nabla\|_{2}^{1-\mathcal{C}_{i}}\quad\text{with}\quad Z=\int\|\nabla\|_{2}^{1-\mathcal{C}_{i}}d\nabla,$ (8)
where $Z$ is the normalization constant. The KL divergence between empirical distribution $p(\nabla)$ and $q^{*}(\nabla)$ is:
$D_{KL}(p\|q^{*})=\mathbb{E}_{p}[\log p(\nabla)]-\mathbb{E}_{p}[\log\|\nabla\|_{2}^{1-\mathcal{C}_{i}}]+\log Z,$ (9)
Substituting the weighted distribution $p_{w}(\nabla)\propto w(\nabla)p_{0}(\nabla)$ and taking variational derivative:
$\frac{\delta D_{KL}}{\delta w}=\frac{(1-\mathcal{C}_{i})\log\|\nabla\|_{2}-\log w}{\mathbb{E}[w]}+\lambda=0,$ (10)
where $\lambda$ is a Lagrange multiplier. Solving this yields the optimal weight:
$w^{*}(\nabla)=\exp\left((1-\mathcal{C}_{i})\log\|\nabla\|_{2}\right).$ (11)
For ease of use and simplification of calculations, we represent $w^{*}(\nabla)$ as $w_{i}\propto(1-\mathcal{C}_{i})$, where $w_{i}$ is the weight corresponding to Gaussian $G_{i}$. Furthermore, we design $w_{i}$ as follows:
$w_{i}=\alpha+\beta\cdot(1-\mathcal{C}_{i})^{p},$ (12)
where $\alpha\in[0,1]$ acts as a baseline suppress factor that suppresses split operations under directional consistency, while $\beta$ serves as an amplification factor to enhance weight responses when directional discrepancies occur. The shape-adjustable hyperparameter $p$ governs the nonlinearity of the weight response curve. As $p$ increases, the curve exhibits progressively steeper gradients, which correspondingly diminish the proportion of Gaussians surpassing the split threshold. This formulation preserves three fundamental mathematical properties: monotonicity, boundedness, and convexity. For visual reference, the response curves corresponding to different $p$ values are presented in Appendix B. The decision measurements are represented as follows:
$\nabla_{split}=\nabla_{\mu_{i}}L\cdot w_{i},\nabla_{clone}=\nabla_{\mu_{i}}L/w_{i}.$ (13)
[…]
Datasets Mip-NeRF360 Tanks&Temples Deep Blending
[…]
Plenoxels fridovich2022plenoxels 0.626 23.08 0.463 2.1GB 0.719 21.08 0.379 2.3GB 0.795 23.06 0.510 2.7GB
INGP muller2022instant 0.671 25.30 0.371 13MB 0.723 21.72 0.330 13MB 0.797 23.62 0.423 13MB
Mip-NeRF360 barron2022mip 0.792 27.69 0.237 8.6MB 0.759 22.22 0.257 8.6MB 0.901 29.40 0.245 8.6MB
3DGS kerbl20233d 0.815 27.21 0.214 734MB 0.841 23.14 0.183 411MB 0.903 29.41 0.243 676MB
Pixel-GS* zhang2024pixelgs 0.823 27.67 0.193 1.2GB 0.853 23.74 0.150 1.05GB 0.896 28.91 0.248 1.1GB
AbsGS-0004 ye2024absgs 0.820 27.49 0.191 728MB 0.853 23.73 0.162 304MB 0.902 29.67 0.236  [^4] MB
AbsGS-0008 ye2024absgs 0.815 27.41 0.211 450MB 0.844 23.54 0.183 202MB 0.903 29.69 0.241 380MB
GDAGS (Ours) 0.821 27.55 0.199 515MB 0.847 23.55 0.176 226MB 0.905 29.70 0.235 388MB
4 Experiments
4.1 Setup
Datasets. Similar to the 3DGS dataset, we selected all 9 scenes introduced in MIP-NERF360 barron2022mip , including indoor and outdoor scenes, Train and Trck scenes in Tanks&Temples hedman2018deep , and drJohnson and playroom scenes in Deep Blending knapitsch2017tanks .
Baselines. The proposed method is benchmarked against three categories of approaches: fast NeRF frameworks (Plenoxels fridovich2022plenoxels , Instant-NGP muller2022instant ), state-of-the-art NeRF-based models (Mip-NeRF360 barron2022mip ), and 3DGS variants including the widely adopted 3DGS kerbl20233d , over-reconstruction-focused AbsGS ye2024absgs , and Pixel-GS zhang2024pixelgs , ensuring comprehensive evaluation across all datasets.
Implementation Details. Our experimental setup follows the original paper parameter settings of 3DGS, with densification performed every 100 iterations, stopping densification after 15k iterations and completing training at 30k iterations. The suppression factor $\alpha$ in the weight function is set to 0.8, the amplification factor $\beta$ is set to the size threshold of 25, and the hyperparameter $p$ is set to 15. All experiments are conducted on a single NVIDIA 4090 GPU with 24GB of graphics memory. We report the view metrics PSNR, SSIM, and LPIPS zhang2018unreasonable , and the memory used to store Gaussian parameters to demonstrate the trade-off between performance and efficiency.
Figure 3: Qualitative comparisons of different methods on scenes from Mip-NeRF360, Tanks&Temples and Deep Blending datasets. Enlarged images are displayed in the bottom right corner.
4.2 Comparison
[…]
Datasets Mip-NeRF360 Tanks&Temples Deep Blending
[…]
3DGS 0.815 27.21 0.214 734MB 0.841 23.14 0.183 411MB 0.903 29.41 0.243 676MB
GDAGS-S 0.819 27.52 0.240 441MB 0.846 23.54 0.178 195MB 0.905 29.67 0.238 328MB
GDAGS-C 0.812 27.46 0.217 615MB 0.847 23.71 0.180 305MB 0.904 29.70 0.240 452MB
GDAGS (Ours) 0.821 27.55 0.199 515MB 0.847 23.55 0.176 226MB 0.905 29.70 0.235 388MB
Quantitative Results. We report the quantitative results of the experiment in the Table 1, which indicate that our method produces comparable results in most cases. It is worth noting that our method has less memory at performance close to or better than SOTA. Compared to Pixel-GS, GDAGS achieves comparable or superior rendering performance with only 20%-50% of the memory consumption. This, combined with its comparison against 3DGS, demonstrates that GDAGS effectively addresses over-densification through nonlinear dynamic weights that suppress redundant Gaussians during densification. By adaptively controlling density, GDAGS enables scene Gaussians to converge toward compact spatial representations. In comparisons under identical memory constraints like AbsGS-0008, GDAGS exhibits full superiority in performance. Even when competing against AbsGS-004 under larger memory budgets, GDAGS maintains comparable results. GDAGS employs direction-aware adaptive dynamic weighting for individual Gaussians, promoting effective densification while optimizing spatial representations.
Qualitative Results. We present novel view synthesis results of GDAGS and baseline methods in Figure 3, demonstrating that our approach achieves high-quality visual fidelity across all scenes while effectively mitigating localized blur and detail loss, such as vegetation under benches, uneven concrete pavements, and textural details on train surfaces. Comprehensive quantitative and qualitative results are presented in the Appendix C.
Figure 4: Performance of different hyperparameters $p$ in multiple datasets.
4.3 Analysis
4.3.1 Ablation analysis
We conduct ablation experiments on the proposed weighting method and trained two additional models: GDAGS-S and GDAGS-C, which represent split gradient weighting and clone gradient weighting, respectively, for experimental comparison with the original 3DGS and complete GDAGS. The experimental results are shown in Table 2. Split gradient weighting can effectively optimize SSIM and LPIPS indicators and reduce memory costs. Clone gradient weighting can improve PSNR indicators but increase memory costs. Integrating split gradient weighting and clone gradient weighting can achieve a balance between performance and efficiency.
4.3.2 Parameter sensitivity analysis
We conduct a parameter sensitivity analysis on the hyperparameter $p$ to reflect the impact of parameters on the model. We set the parameters to 10, 15, 20, 25, and 30, respectively, and the specific data is shown in Figure 4. In the previous analysis, we explained that the hyperparameter $p$ controls the number of Gaussians selected for densification in adaptive density control, thereby indirectly controlling the total number of Gaussians. The increase of hyperparameter $p$ decreases the number of Gaussians used for densification, thereby reducing memory and lowering rendering performance. The reduction of hyperparameter $p$ increases the number of Gaussians used for densification, thereby increasing memory and improving rendering performance.
Figure 5: Visualization of different densification methods during the training process in bicycle sense.
4.3.3 Densification process analysis
We quantified the Gaussian split/clone counts per densification round and the total Gaussian population dynamics during training (see Figure 5). The experimental results demonstrate distinct behavioral patterns: In cloning operations, all three methods exhibit consistent trends, with 3DGS achieving the highest clone count due to its standard threshold. AbsGS, employing twice the cloning threshold of 3DGS, yields the lowest clone numbers, whereas GDAGS suppresses directionally inconsistent clones through gradient alignment constraints, resulting in intermediate cloning activity. GDAGS further ensures localized replenishment of directionally consistent clones while mitigating redundant accumulation of geometrically conflicting clones. For splitting operations, AbsGS generates excessive splits by indiscriminately utilizing gradient magnitudes, whereas 3DGS underperforms due to gradient direction conflicts. GDAGS effectively filters splittable Gaussians through adaptive criteria. Regarding total Gaussian counts, GDAGS exhibits a more stable trajectory and faster convergence compared to 3DGS and AbsGS, validating its optimized densification strategy in balancing geometric fidelity and computational efficiency.
5 Conclusion
This paper proposes GDAGS, a dynamic adaptive Gaussian splatting framework that addresses critical challenges in 3D scene reconstruction, including over-densification, over-reconstruction, and memory inefficiency. By introducing the directional consistency metric (GCR) and nonlinear dynamic weighting, GDAGS enables direction-aware densification control: GCR discriminates between directionally consistent/divergent Gaussians, while dynamic weights independently regulate densification at the per-Gaussian level. This dual mechanism effectively suppresses redundant Gaussian proliferation, eliminates outlier distributions, and drives compact spatial representations without relying on heuristic threshold tuning.
Appendix A Algorithm and weighted response curve
Algorithm 1 Gradient-Direction-Aware Adaptive Densification
1:Gaussian set $\mathcal{G}$, training views $\mathcal{V}$, initial thresholds $\tau$
2:Optimized Gaussian set $\mathcal{G}^{\prime}$
3:Initialize gradient buffers $\{\nabla_{i,v}\}\leftarrow 0$
4:for each iteration $t=1$ to $T$ do
5:  for each view $v\in\mathcal{V}$ do
6:   Render $\mathcal{G}$ via differentiable splatting
7:   Compute 2D gradient map $\sum_{pixel}\nabla_{i,pixel}^{v}$ and $\sum_{pixel}\|\nabla_{i,p}^{v}\|_{2}$
8:   Accumulate $\nabla_{i}^{v}$ for visible Gaussians
9:  end for
10:  for each Gaussian $g_{i}\in\mathcal{G}$ do
11:   Compute $\mathcal{C}_{i}$ via Eq. (5)
12:   Calculate $w_{i}$ using Eq. (12)
13:  end for
14:  Compute $\nabla_{split}$, $\nabla_{clone}$
15:  Split if $\nabla_{split}>\tau_{p}~\&~\Sigma_{3D}^{i}>\tau_{s}$
16:  Clone if $\nabla_{clone}>\tau_{p}~\&~\Sigma_{3D}^{i}\leq\tau_{s}$
17:end for
Appendix B Gaussian densification
Figure 6: Corresponding curves of weights for different hyperparameters pp.
Figure 7: Comparison of more detailed methods in the Gaussian densification process.
Figure 8: Visualization of different densification methods during the training process in flowers sense.
Appendix C Detailed results
The detailed performance metrics across a total of 13 different scenes from the three datasets are shown in Tables 34578, and 9 while efficiency metrics are presented in Tables 6 and 10. Table 11 presents the statistical error analysis. Render quality comparisons are demonstrated in Figures 9 and 10, where significant differences are highlighted with red boxes and enlarged in the bottom-right corners. Figure 11 compares over-densification phenomena, with FPS values and the number of Gaussians annotated in the top-left corners. The comparison between nonlinear weights and linear weights is shown in the Table 12.
Table 3: Per-scene quantitative results (SSIM) from the Mip-NeRF360.
bicycle flowers garden stump treehill room counter kitchen bonsai Plenoxels 0.496 0.431 0.606 0.523 0.509 0.841 0.759 0.648 0.814 INGP-Base 0.491 0.450 0.649 0.574 0.518 0.855 0.798 0.818 0.890 INGP-Big 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906 Mip-NeRF360 0.685 0.583 0.813 0.744 0.632 0.913 0.894 0.920 0.941 3D-GS 0.771 0.605 0.868 0.775 0.638 0.914 0.905 0.922 0.938 AbsGS-0004 0.782 0.613 0.870 0.784 0.626 0.920 0.908 0.929 0.944 Pixel-GS 0.766 0.627 0.862 0.784 0.638 0.929 0.921 0.936 0.951 GDAGS 0.776 0.615 0.861 0.786 0.624 0.930 0.915 0.933 0.948
Table 4: Per-scene quantitative results (PSNR) from the Mip-NeRF360.
bicycle flowers garden stump treehill room counter kitchen bonsai Plenoxels 21.912 20.097 23.495 20.661 22.248 27.594 23.624 23.420 24.669 INGP-Base 22.193 20.348 24.599 23.626 22.364 29.269 26.439 28.548 30.337 INGP-Big 22.171 20.652 25.069 23.466 22.373 29.690 26.691 29.479 30.685 Mip-NeRF360 24.37 21.73 26.98 26.40 22.87 31.63 29.55 32.23 33.46 3D-GS 25.246 21.520 27.410 26.550 22.490 30.632 28.700 30.317 31.980 AbsGS-0004 25.373 21.298 27.579 26.766 22.074 31.582 28.968 31.774 32.283 Pixel-GS 25.263 21.677 27.393 26.938 22.343 31.952 29.270 31.950 32.687 GDAGS 25.309 21.408 27.465 26.960 22.034 31.820 29.015 31.951 32.277
Table 5: Per-scene quantitative results (LPIPS) from the Mip-NeRF360.
bicycle flowers garden stump treehill room counter kitchen bonsai Plenoxels 0.506 0.521 0.386 0.503 0.540 0.4186 0.441 0.447 0.398 INGP-Base 0.487 0.481 0.312 0.450 0.489 0.301 0.342 0.254 0.227 INGP-Big 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205 Mip-NeRF360 0.301 0.344 0.170 0.261 0.339 0.211 0.204 0.127 0.176 3D-GS 0.205 0.336 0.103 0.210 0.317 0.220 0.204 0.129 0.205 AbsGS-0004 0.186 0.295 0.104 0.202 0.297 0.216 0.198 0.124 0.194 Pixel-GS 0.206 0.288 0.111 0.208 0.301 0.188 0.166 0.109 0.167 GDAGS 0.195 0.297 0.117 0.215 0.310 0.186 0.182 0.116 0.172
Table 6: Per-scene memory consumption (MB) from the Mip-NeRF360.
bicycle flowers garden stump treehill room counter kitchen bonsai 3D-GS* 1508 813 1073 1100 1018 397 257 412 259 Pixel-GS 2040 1660 1980 1480 1690 584 592 721 487 AbsGS-0004 1156 751 932 1021 962 234 192 249 234 GDAGS 1030 635 841 683 744 223 146 197 144
Table 7: Per-scene quantitative results (SSIM) from the Tanks & Temples and Deep Blending.
Truck Train Dr Johnson Playroom Plenoxels 0.774 0.663 0.787 0.802 INGP-Base 0.779 0.666 0.839 0.754 INGP-Big 0.800 0.689 0.854 0.779 Mip-NeRF360 0.857 0.660 0.901 0.900 3D-GS 0.879 0.802 0.899 0.906 AbsGS-0004 0.886 0.820 0.900 0.907 Pixel-GS 0.886 0.826 0.887 0.905 GDAGS 0.884 0.811 0.901 0.909
Table 8: Per-scene quantitative results (PSNR) from the Tanks & Temples and Deep Blending.
Truck Train Dr Johnson Playroom Plenoxels 23.221 23.221 23.142 22.980 INGP-Base 23.260 20.170 27.750 19.483 INGP-Big 23.383 20.456 28.257 21.665 Mip-NeRF360 24.912 19.523 29.140 29.657 3D-GS 25.187 21.097 28.766 30.044 AbsGS-0004 25.702 22.010 28.930 29.967 Pixel-GS 25.467 22.028 28.128 29.700 GDAGS 25.560 21.557 29.232 30.166
Table 9: Per-scene quantitative results (LPIPS) from the Tanks & Temples and Deep Blending.
Truck Train Dr Johnson Playroom Plenoxels 0.335 0.422 0.521 0.499 INGP-Base 0.274 0.386 0.381 0.465 INGP-Big 0.249 0.360 0.352 0.428 Mip-NeRF360 0.159 0.354 0.237 0.252 3D-GS 0.148 0.218 0.244 0.241 AbsGS-0004 0.131 0.193 0.240 0.232 Pixel-GS 0.120 0.178 0.255 0.241 GDAGS 0.143 0.208 0.238 0.233
Table 10: Per-scene memory consumption (MB) from the Tanks & Temples and Deep Blending.
Truck Train Dr Johnson Playroom 3D-GS* 530 218 744 504 AbsGS-0004 419 210 457 316 Pixel-GS 1200 904 1280 881 GDAGS 277 175 416 280
Table 11: Error analysis.
[…]
Table 12: The LPIPS comparison between nonlinear weights and linear weights. Linear weight set to $2-\mathcal{C}_{i}$.
playroom Train bicycle flowers nonlinear weights 0.233 0.208 0.195 0.297 linear weights 0.238 0.210 0.218 0.340
Figure 9: Supplementary qualitative analysis results.
Figure 10: Supplementary qualitative analysis results.
Figure 11: Comparing over-densification phenomena.
Appendix D Limitations of GDAGS
While GDAGS demonstrates significant improvements in adaptive densification control and memory efficiency, several limitations remain:
  • Fixed Hyperparameters for Weighting Schemes: the hyperparameter $p$, which balances performance and efficiency, is fixed at 15 by default. While this value works well empirically, its optimal setting may vary across datasets or hardware configurations, requiring manual tuning for specific use cases.
  • Challenges in extreme sparsity: in highly sparse regions with minimal gradient activity, GCR may struggle to distinguish between true outliers and under-reconstructed areas, leading to potential under-densification or over-suppression of Gaussians.